OpenStudy (anonymous):

Find three consecutive integers such that twice the greatest integer is 2 less than 3 times the least integer.

OpenStudy (chihiroasleaf):

let the integers be a-1, a, a+1 then you have 2(a+1) + 2 = 3 ( a- 1) solve the latter equation, then you'll find the value of a

OpenStudy (anonymous):

The three consecutive integers are x, x+1, and x+2. Then 2(x+2)=3(x)-2. Solve the equation and then substitute to find the three numbers